OpenStudy (anonymous):

f(x)=1/4(x+6)^2+3 What is the vertex

OpenStudy (anonymous):

vertex= ( -6,3)

OpenStudy (anonymous):

I got(-12x,0)

OpenStudy (anonymous):

how did you get your answer

OpenStudy (anonymous):

f(x)=1/4(x+6)^2+3 y=3 (x+6)=0 x=-6

OpenStudy (anonymous):

let me work on paper and get back